OpenStudy (kakarot):

solve for y 5 = 2y + 3

OpenStudy (anonymous):

5= 2y +3 Subtract 3 on both sides 5 - 3 = 2y + 3 - 3 2 = 2y Divide 2 on both sides 2/2 = 2y/2 Simplify 1 = y Answer is y = 1

OpenStudy (kakarot):

ok thanks